# Revisiting Multimodal KV Cache Compression: A Frequency-Domain-Guided Outlier-KV-Aware Approach

### v1

@@ 1 Introduction
Figure 1: Frequency-domain energy distribution of the KV matrices. We observe that the frequency-domain energy of KV matrices is predominantly concentrated at low frequency, with high frequency components occupying a relatively small proportion. The red dashed box in the figure highlights the low-frequency concentration phenomenon. Experiments are conducted on MileBench [song2024milebench] with Qwen2.5-VL-7B.
@@
Multimodal large language models (MLLMs) [achiam2023gpt, zhu2024multilingual, chen2024internvl, li2024survey, liu2024improved, wang2024qwen2] exhibit strong multimodal perception and reasoning, yet their practical deployment is hindered by high computation and memory costs, driven primarily by the substantial volume of visual tokens at inference time [jin2024efficient, huang2024dynamic]. To address this, many techniques, such as token pruning [chen2024image, he2025zipvl, tan2025tokencarve, zhang2024sparsevlm], model quantization [yu2025mquant, xie2024advancing], and KV Cache optimizations [adnan2024keyformer, hooper2024kvquant, li2024survey1], have been proposed to accelerate MLLM inference. Among them, KV Cache improves decoding efficiency by storing the Keys and Values of previously seen tokens, avoiding repeated prefix computation during decoding, thereby garnering significant attention [shi2024keep, ge2023model]. Despite achieving impressive progress, as context length grows (e.g., multi-image, high-resolution and video settings [song2024milebench, meng2024mmiu, hrbench, tu2025favor]), the length of multimodal KV Cache increases accordingly, leading to substantial GPU memory overhead and a sharp slowdown during the decoding stage.
@@
Recently, several works have attempted to prune redundant portions of long-context KV Cache in MLLMs to further reduce memory usage and speed up decoding. LOOK-M [wan2024look] observes extensive redundancy in visual KVs within long multimodal contexts and compresses the KV Cache by screening and merging visual KV pairs with low attention score during the prefill stage. MEDA [wan2025meda] uses cross-modal attention entropy to allocate an appropriate KV Cache size per MLLM layer. However, these approaches all rely on attention score to shrink the KV Cache, which raises several issues: (1) efficient attention kernels (e.g., FlashAttention [dao2022flashattention]) typically do not explicitly output the full attention score, and recomputing attention score introduces additional overhead; (2) since attention score are obtained via Query–Key dot products, directly compressing KV Cache via attention score ignores the information contribution of the Value matrix [senguptavalue]. This naturally leads to the question: how can we perform efficient multimodal KV Cache compression without relying on attention score?
@@
Figure 2: The impact on model performance of prioritizing the removal of KV pairs under different removal ratios. We find discarding KV pairs with larger deviations leads to faster performance drop than random removal or removing small-deviation pairs. Experiments are conducted on MileBench [song2024milebench] with Qwen2.5-VL-7B.
@@
In this work, we revisit the problem from the perspective of the data distribution of the visual KV matrices themselves. Inspired by the common use of frequency-domain analysis in image processing [wallace1991jpeg], we transform visual KV matrices into the frequency domain and analyze their energy distribution (i.e., the average power spectrum). As shown in Fig. 1, we observe that the frequency-domain energy of KV matrices is predominantly concentrated at low frequencies, with high-frequency components occupying a relatively small proportion. Further, noting that removing outliers in model quantization often causes steep performance drops [dettmers2022gpt3, xiao2023smoothquant, frantar2022gptq], we hypothesize that a similar outlier phenomenon may exist for KV Cache compression. To investigate, we apply a low-pass filter to obtain smoothed and dominant KV matrices (termed as Base KV), then measure the deviation of the original KV matrices from Base KV and remove KV pairs with different deviations. We find that, discarding KV pairs with larger deviations leads to faster performance drop than random removal or removing small-deviation pairs (Fig. 2), whereas discarding small-deviation KV pairs has smaller impact on performance. These results indicate KV pairs that differ substantially from the Base KV are more likely to carry features critical to model inference, which we define as Outlier KVs.
@@
Building on this novel finding, we propose FlashCache, a frequency-domain-guided Outlier-KV-aware compression approach with two core components. First, we introduce an Outlier KV Recognition Module that efficiently selects Outlier KVs in each layer. Concretely, we apply Discrete Cosine Transform (DCT) [ahmed2006discrete] to map the KV matrices to the frequency domain and retain only the dominant low-frequency components; we then perform the inverse DCT to return to the time domain and obtain the smoothed Base KV; after that, we compute the mean squared error between each original KV pair and its Base KV counterpart to select Outlier KVs. Second, we further observe differences in the low-frequency concentration phenomenon across each layer of the model in Fig. 4, and propose a frequency-guided Dynamic Budget Allocation Module to dynamically determine the KV Cache size per layer. For each layer, we analyze the KV matrices in the frequency domain and compute the ratio of outlier-information energy to the total KV energy; we normalize these ratios into weights and allocate distinct per-layer retention quotas under a global cache budget. Across multiple models and datasets, FlashCache reduces KV usage and accelerates decoding under the same context length while preserving model performance.
@@ 2.1 Multimodal Large Language Models
Recently, multimodal large language models (MLLMs) have demonstrated powerful image-text understanding capabilities [liu2024mmbench, chen2024internvl, zhang2024long, li2024mini]. LLaVA [liu2024improved] aligns a visual encoder with a large language model via visual instruction tuning to enable general image-to-text understanding and multi-turn reasoning. InternVL [chen2024internvl] couples a strong vision backbone with a language model, introducing multi-granularity visual tokens and high-resolution modeling for improved detail perception and cross-task transfer. QwenVL [wang2024qwen2] extends the Qwen LLM with vision, supporting multi-image inputs, OCR, and referring expression grounding to strengthen open-domain multimodal reasoning. However, as context length increases (e.g., multi-image, high-resolution and video settings [song2024milebench, meng2024mmiu, hrbench, tu2025favor]), the multimodal KV Cache of MLLMs grows accordingly, leading to excessive GPU memory consumption and slower decoding speeds. Therefore, reducing the size of the KV Cache generated during long-context multimodal inference to accelerate decoding without compromising performance is both a necessary and valuable challenge.
@@ 2.2 Visual Compression for MLLMs
In long-context scenarios, visual information experiences explosive growth. On one hand, in terms of volume, visual tokens often exceeds textual tokens by hundreds or even thousands of times [chen2024image]. On the other hand, visual signals are inherently sparser in information compared to texts produced by humans [marr2010vision]. Therefore, it is necessary to compress visual information during inference process of long-context MLLMs. Existing methods for compressing visual information in long-context MLLMs mainly include: visual token compression [chen2024image, he2025zipvl, tan2025tokencarve, zhang2024sparsevlm, shang2025llava], quantization [yu2025mquant, xie2024advancing], and multimodal KV Cache compression [li2024survey, he2024zipvl].
Among these, the KV Cache compression method has garnered significant attention for its ability to maintain performance while achieving high compression ratios, thereby reducing memory usage and accelerating decoding. LOOK-M [wan2024look] compresses the KV Cache by screening and merging visual KV pairs with low attention score once during the prefill stage. MEDA [wan2025meda] uses cross-modal attention entropy to allocate an appropriate KV Cache size per MLLM layer. However, these methods are not compatible with established efficient attention kernels ( i.e., FlashAttention [dao2022flashattention]) and overlook the contribution of value vectors, which directly influence the attention output. Therefore, we propose FlashCache, a frequency-domain-guided, training-free hierarchical KV compression framework for efficient multimodal long-context inference.
@@ 3.2.1 Outlier KV Recognition Module
where $l$ represents model layer numbers, we find that their frequency domain energy intensity is concentrated at low frequency as shown in Fig. 1. In natural signals, a large portion of the total energy is often concentrated in lower-frequency components, which correspond to the smooth, structural information. According to Parseval’s theorem [iwasaki2019deriving], the energy of a signal is invariant between spatial and frequency domains. Therefore, we argue that retaining the principal components of the KV matrices in the frequency domain preserves the fundamental trends in KV representation. We therefore perform the low-pass filter of the original Key and Value tensors to obtain a smooth KV representation that is defined as Base KV, which is able to characterize the primary trends in the KV matrices.
@@
Figure 4: Frequency-domain energy distribution of KV matrices across different layers. We observe significant differences in the frequency-domain energy distribution of KV matrices across different layers. The red dashed box indicates the energy distributions of KV matrices of different layers show different modes. Experiments are conducted on MileBench [song2024milebench] with Qwen2.5-VL-7B.
@@
StreamingLLM [xiao2023efficient] 46.97 65.04 9.22 12.17 55.59 67.51 9.69 14.00 56.75 63.15 3.75 19.33
H2O [zhang2023h2o] 46.97 65.29 19.69 12.33 55.59 68.60 12.66 14.67 57.01 64.73 14.32 18.00
SnapKV [li2024snapkv] 46.97 65.27 19.37 12.17 55.59 68.27 13.59 15.33 57.07 65.61 22.34 18.83
LOOK-M [wan2024look] 46.97 65.23 18.60 11.83 55.55 67.50 11.88 11.83 51.03 53.37 5.78 13.00
MEDA [wan2025meda] 46.91 65.08 19.97 12.33 55.59 68.13 9.07 14.50 51.86 54.01 5.16 14.50
@@ 3.2.2 Dynamic Budget Allocation Module
Previous studies [zhang2024sparsevlm, cai2024pyramidkv, li2024survey] have found that different layers of MLLMs exhibit varying degrees of information redundancy. Allocating distinct KV Cache budgets to each layer maximizes the reduction of information loss caused by compression. At the same time, we observe that the low-frequency concentration in the frequency domain of the KV matrices across different layers of the model is not consistent, as shown in Fig. 4. In particular, different transformer layers exhibit varying proportions of outlier energy: in the frequency domain, the key-value matrices of certain layers are predominantly concentrated in low-frequency principal components, while other layers carry relatively more high-frequency outlier energy. This cross-layer disparity suggests that a uniform KV retention ratio for all layers would be suboptimal. Based on this, we further propose a frequency-domain guided Dynamic Budget Allocation Module to dynamically determine the size of the KV Cache for each layer.
@@ 4.1 Experimental Setup
Datasets and models. To validate the effectiveness of our approach, we evaluate FlashCache on diverse benchmarks spanning multiple tasks: three multi-image understanding benchmarks (MileBench [song2024milebench], MUIRBench [wang2024muirbench] and MMMU [yue2023mmmu]), two high-resolution benchmarks (V* [vstar] and HR-Bench [hrbench]) and one video benchmark (FAVOR-Bench [tu2025favor]). We verify FlashCache on three MLLMs: LLaVA-OneVision-1.5-8B-Instruct [LLaVAOneVision15] and Qwen2.5-VL-7B/32B-Instruct [wang2024qwen2].
Implementation details. All experiments are conducted on a single NVIDIA H200 (141GB) GPU with FlashAttention [dao2022flashattention]. We compare FlashCache against state-of-the-art KV cache eviction methods: StreamingLLM [xiao2023efficient], H2O [zhang2023h2o], SnapKV [li2024snapkv], LOOK-M [wan2024look], and MEDA [wan2025meda]. We utilize NVIDIA’s CuPy operator library to accelerate the DCT operation within FlashCache.
@@ 4.2 Experiments on Multi-Images Understanding Benchmarks
StreamingLLM [xiao2023efficient] 37.58 37.58 50.0 50.0
H2O [zhang2023h2o] OOM OOM 53.18 51.18
SnapKV [li2024snapkv] OOM 44.42 53.18 52.18
LOOK-M [wan2024look] OOM OOM 53.18 52.18
MEDA [wan2025meda] OOM OOM 53.18 51.36
@@
StreamingLLM [xiao2023efficient] 78.77 78.65 71.25 71.25
H2O [zhang2023h2o] 78.77 78.77 70.75 71.12
SnapKV [li2024snapkv] 79.56 78.89 70.88 71.12
LOOK-M [wan2024look] 78.31 77.78 70.0 70.25
MEDA [wan2025meda] 78.84 77.78 OOM OOM
@@ 4.3 Experiments on High-Resolution Benchmarks
StreamingLLM [xiao2023efficient] 30.94 39.13 39.05 40.5 31.35 28.12 35.57
H2O [zhang2023h2o] 32.16 38.22 39.05 40.66 31.35 23.44 35.78
SnapKV [li2024snapkv] 28.9 39.71 40.91 39.83 33.49 39.06 35.67
LOOK-M [wan2024look] 30.38 40.23 39.53 43.4 32.56 25.0 36.25
MEDA [wan2025meda] 26.17 36.28 37.61 39.75 31.91 21.88 33.11
@@
StreamingLLM [xiao2023efficient] 20.52 29.4 32.55 31.29 28.0 12.5 27.14
H2O [zhang2023h2o] 21.35 28.68 32.19 31.45 28.56 12.5 27.3
SnapKV [li2024snapkv] 24.19 32.25 34.66 34.69 29.02 15.62 29.95
LOOK-M [wan2024look] 20.29 30.11 32.49 33.61 31.81 12.5 28.03
MEDA [wan2025meda] 15.17 24.72 25.99 29.13 27.53 12.5 22.83
@@ 4.6 Ablation Study
Ablation Study on the low-pass filter cut-off factor $\gamma$. In this section, we first examine the impact of $\gamma$ on FlashCache, which is used as a control parameter for the cutoff frequency of the low-pass filter in Sec. 2.2. Specifically, we employ Qwen2.5-VL-7B for the experiment, with the KV retention ratio set to 0.2 and the range of values for $\gamma$ from 0 to 1. To facilitate comparison of differences, we conduct experiments on the ImageNeedleInAHaystack and GPR1200 datasets (both included in MileBench). As shown in Tab. 5, we observe that as $\gamma$ increases, the model’s performance gradually declines. Optimal model performance is achieved when $\gamma$ falls within the range of 0.1 to 0.3. Since the KV matrix is predominantly concentrated in low-frequency ranges, excessive $\gamma$ values prevent the low-pass filter from effectively extracting the primary representation of the KV matrix (Base KV). This failure to identify the Outlier KV consequently leads to a decline in model performance. This phenomenon is consistent with our core findings.
@@
Method H2O [zhang2023h2o] SnapKV [li2024snapkv] LOOK-M [wan2024look] MEDA [wan2025meda] FlashCache

### v2

@@ 1 Introduction
Figure 1: Frequency-domain energy distribution of the KV matrices. We observe that the frequency-domain energy of KV matrices is predominantly concentrated at low frequency, with high frequency components occupying a relatively small proportion. The red dashed box in the figure highlights the low-frequency concentration phenomenon. Experiments are conducted on MileBench [29] with Qwen2.5-VL-7B.
@@
Multimodal large language models (MLLMs) [1, 50, 7, 19, 22, 37] exhibit strong multimodal perception and reasoning, yet their practical deployment is hindered by high computation and memory costs, driven primarily by the substantial volume of visual tokens at inference time [17, 15, 43]. To address this, many techniques, such as token pruning [6, 12, 31, 48], model quantization [45, 42], and KV Cache optimizations [2, 13, 18], have been proposed to accelerate MLLM inference. Among them, KV Cache improves decoding efficiency by storing the Keys and Values of previously seen tokens, avoiding repeated prefix computation during decoding, thereby garnering significant attention [28, 11]. Despite achieving impressive progress, as context length grows (e.g., multi-image, high-resolution and video settings [29, 25, 38, 32]), the length of multimodal KV Cache increases accordingly, leading to substantial GPU memory overhead and a sharp slowdown during the decoding stage.
@@
Recently, several works have attempted to prune redundant portions of long-context KV Cache in MLLMs to further reduce memory usage and speed up decoding. LOOK-M [35] observes extensive redundancy in visual KVs within long multimodal contexts and compresses the KV Cache by screening and merging visual KV pairs with low attention score during the prefill stage. MEDA [34] uses cross-modal attention entropy to allocate an appropriate KV Cache size per MLLM layer. However, these approaches all rely on attention score to shrink the KV Cache, which raises several issues: (1) efficient attention kernels (e.g., FlashAttention [8]) typically do not explicitly output the full attention score, and recomputing attention score introduces additional overhead; (2) since attention score are obtained via Query–Key dot products, directly compressing KV Cache via attention score ignores the information contribution of the Value matrix [26]. This naturally leads to the question: how can we perform efficient multimodal KV Cache compression without relying on attention score?
@@
Figure 2: The impact on model performance of prioritizing the removal of KV pairs under different removal ratios. We find discarding KV pairs with larger deviations leads to faster performance drop than random removal or removing small-deviation pairs. Experiments are conducted on MileBench with Qwen2.5-VL-7B.
@@
In this work, we revisit the problem from the perspective of the data distribution of the visual KV matrices themselves. Inspired by the common use of frequency-domain analysis in image processing [33], we transform visual KV matrices into the frequency domain and analyze their energy distribution (i.e., the average power spectrum). As shown in Fig. 1, we observe that the frequency-domain energy of KV matrices is predominantly concentrated at low frequencies, with high-frequency components occupying a relatively small proportion. Further, noting that removing outliers in model quantization often causes steep performance drops [9, 40, 10], we hypothesize that a similar outlier phenomenon may exist for KV Cache compression. To investigate, we apply a low-pass filter to obtain smoothed and dominant KV matrices (termed as Base KV), then measure the deviation of the original KV matrices from Base KV and remove KV pairs with different deviations. We find that, discarding KV pairs with larger deviations leads to faster performance drop than random removal or removing small-deviation pairs (Fig. 2), whereas discarding small-deviation KV pairs has smaller impact on performance. These results indicate KV pairs that differ substantially from the Base KV are more likely to carry features critical to model inference, which we define as Outlier KVs.
@@
Building on this novel finding, we propose FlashCache, a frequency-domain-guided Outlier-KV-aware compression approach with two core components. First, we introduce an Outlier KV Recognition Module that efficiently selects Outlier KVs in each layer. Concretely, we apply Discrete Cosine Transform (DCT) [3] to map the KV matrices to the frequency domain and retain only the dominant low-frequency components; we then perform the inverse DCT to return to the time domain and obtain the smoothed Base KV; after that, we compute the mean squared error between each original KV pair and its Base KV counterpart to select Outlier KVs. Second, we further observe differences in the low-frequency concentration phenomenon across each layer of the model in Fig. 4, and propose a frequency-guided Dynamic Budget Allocation Module to dynamically determine the KV Cache size per layer. For each layer, we analyze the KV matrices in the frequency domain and compute the ratio of outlier-information energy to the total KV energy; we normalize these ratios into weights and allocate distinct per-layer retention quotas under a global cache budget. Across multiple models and datasets, FlashCache reduces KV usage and accelerates decoding under the same context length while preserving model performance.
@@ 2.1 Multimodal Large Language Models
Recently, multimodal large language models (MLLMs) have demonstrated powerful image-text understanding capabilities [23, 7, 47, 20, 44, 14]. LLaVA [22] aligns a visual encoder with a large language model via visual instruction tuning to enable general image-to-text understanding and multi-turn reasoning. InternVL [7] couples a strong vision backbone with a language model, introducing multi-granularity visual tokens and high-resolution modeling for improved detail perception and cross-task transfer. QwenVL [37] extends the Qwen LLM with vision, supporting multi-image inputs, OCR, and referring expression grounding to strengthen open-domain multimodal reasoning. However, as context length increases (e.g., multi-image, high-resolution and video settings [29, 25, 38, 32]), the multimodal KV Cache of MLLMs grows accordingly, leading to excessive GPU memory consumption and slower decoding speeds. Therefore, reducing the size of the KV Cache generated during long-context multimodal inference to accelerate decoding without compromising performance is both a necessary and valuable challenge.
@@ 2.2 Visual Compression for MLLMs
In long-context scenarios, visual information experiences explosive growth. On one hand, in terms of volume, visual tokens often exceeds textual tokens by hundreds or even thousands of times [6]. On the other hand, visual signals are inherently sparser in information compared to texts produced by humans [24]. Therefore, it is necessary to compress visual information during inference process of long-context MLLMs. Existing methods for compressing visual information in long-context MLLMs mainly include: visual token compression [6, 31, 48, 27, 30], quantization [45, 42], and multimodal KV Cache compression [19, 12].
Among these, the KV Cache compression method has garnered significant attention for its ability to maintain performance while achieving high compression ratios, thereby reducing memory usage and accelerating decoding. LOOK-M [35] compresses the KV Cache by screening and merging visual KV pairs with low attention score once during the prefill stage. MEDA [34] uses cross-modal attention entropy to allocate an appropriate KV Cache size per MLLM layer. However, these methods are not compatible with established efficient attention kernels ( i.e., FlashAttention [8]) and overlook the contribution of value vectors, which directly influence the attention output. Therefore, we propose FlashCache, a frequency-domain-guided, training-free hierarchical KV compression framework for efficient multimodal long-context inference.
@@ 3.2.1 Outlier KV Recognition Module
where $l$ represents model layer numbers, we find that their frequency domain energy intensity is concentrated at low frequency as shown in Fig. 1. In natural signals, a large portion of the total energy is often concentrated in lower-frequency components, which correspond to the smooth, structural information. According to Parseval’s theorem [16], the energy of a signal is invariant between spatial and frequency domains. Therefore, we argue that retaining the principal components of the KV matrices in the frequency domain preserves the fundamental trends in KV representation. We therefore perform the low-pass filter of the original Key and Value tensors to obtain a smooth KV representation that is defined as Base KV, which is able to characterize the primary trends in the KV matrices.
@@
Figure 4: Frequency-domain energy distribution of KV matrices across different layers. We observe significant differences in the frequency-domain energy distribution of KV matrices across different layers. The red dashed box indicates the energy distributions of KV matrices of different layers show different modes. Experiments are conducted on MileBench [29] with Qwen2.5-VL-7B.
@@
StreamingLLM [41] 46.97 65.04 9.22 12.17 55.59 67.51 9.69 14.00 56.75 63.15 3.75 19.33
H2O [49] 46.97 65.29 19.69 12.33 55.59 68.60 12.66 14.67 57.01 64.73 14.32 18.00
SnapKV [21] 46.97 65.27 19.37 12.17 55.59 68.27 13.59 15.33 57.07 65.61 22.34 18.83
LOOK-M [35] 46.97 65.23 18.60 11.83 55.55 67.50 11.88 11.83 51.03 53.37 5.78 13.00
MEDA [34] 46.91 65.08 19.97 12.33 55.59 68.13 9.07 14.50 51.86 54.01 5.16 14.50
@@ 3.2.2 Dynamic Budget Allocation Module
Previous studies [48, 5, 19] have found that different layers of MLLMs exhibit varying degrees of information redundancy. Allocating distinct KV Cache budgets to each layer maximizes the reduction of information loss caused by compression. At the same time, we observe that the low-frequency concentration in the frequency domain of the KV matrices across different layers of the model is not consistent, as shown in Fig. 4. In particular, different transformer layers exhibit varying proportions of outlier energy: in the frequency domain, the key-value matrices of certain layers are predominantly concentrated in low-frequency principal components, while other layers carry relatively more high-frequency outlier energy. This cross-layer disparity suggests that a uniform KV retention ratio for all layers would be suboptimal. Based on this, we further propose a frequency-domain guided Dynamic Budget Allocation Module to dynamically determine the size of the KV Cache for each layer.
@@ 4.1 Experimental Setup
Datasets and models. To validate the effectiveness of our approach, we evaluate FlashCache on diverse benchmarks spanning multiple tasks: three multi-image understanding benchmarks (MileBench [29], MUIRBench [36] and MMMU [46]), two high-resolution benchmarks (V* [39] and HR-Bench [38]) and one video benchmark (FAVOR-Bench [32]). We verify FlashCache on three MLLMs: LLaVA-OneVision-1.5-8B-Instruct [4] and Qwen2.5-VL-7B/32B-Instruct [37].
Implementation details. All experiments are conducted on a single NVIDIA H200 (141GB) GPU with FlashAttention [8]. We compare FlashCache against state-of-the-art KV cache eviction methods: StreamingLLM [41], H2O [49], SnapKV [21], LOOK-M [35], and MEDA [34]. We utilize NVIDIA’s CuPy operator library to accelerate the DCT operation within FlashCache.
@@ 4.2 Experiments on Multi-Images Understanding Benchmarks
StreamingLLM [41] 37.58 37.58 50.0 50.0
H2O [49] OOM OOM 53.18 51.18
SnapKV [21] OOM 44.42 53.18 52.18
LOOK-M [35] OOM OOM 53.18 52.18
MEDA [34] OOM OOM 53.18 51.36
@@
StreamingLLM [41] 78.77 78.65 71.25 71.25
H2O [49] 78.77 78.77 70.75 71.12
SnapKV [21] 79.56 78.89 70.88 71.12
LOOK-M [35] 78.31 77.78 70.0 70.25
MEDA [34] 78.84 77.78 OOM OOM
@@ 4.3 Experiments on High-Resolution Benchmarks
StreamingLLM [41] 30.94 39.13 39.05 40.5 31.35 28.12 35.57
H2O [49] 32.16 38.22 39.05 40.66 31.35 23.44 35.78
SnapKV [21] 28.9 39.71 40.91 39.83 33.49 39.06 35.67
LOOK-M [35] 30.38 40.23 39.53 43.4 32.56 25.0 36.25
MEDA [34] 26.17 36.28 37.61 39.75 31.91 21.88 33.11
@@
StreamingLLM [41] 20.52 29.4 32.55 31.29 28.0 12.5 27.14
H2O [49] 21.35 28.68 32.19 31.45 28.56 12.5 27.3
SnapKV [21] 24.19 32.25 34.66 34.69 29.02 15.62 29.95
LOOK-M [35] 20.29 30.11 32.49 33.61 31.81 12.5 28.03
MEDA [34] 15.17 24.72 25.99 29.13 27.53 12.5 22.83
@@ 4.6 Ablation Study
Ablation Study on the low-pass filter cut-off factor $\gamma$. In this section, we first examine the impact of $\gamma$ on FlashCache, which is used as a control parameter for the cutoff frequency of the low-pass filter in Sec. 2.2. Specifically, we employ Qwen2.5-VL-7B for the experiment, with the KV retention ratio set to 0.2 and the range of values for $\gamma$ from 0 to 1. To facilitate comparison of differences, we conduct experiments on the ImageNeedleInAHaystack and GPR1200 datasets (both included in MileBench). As shown in Tab. 5, we observe that as $\gamma$ increases, the model’s performance gradually declines. Optimal model performance is achieved when $\gamma$ falls within the range of 0.1 to 0.2. Since the KV matrix is predominantly concentrated in low-frequency ranges, excessive $\gamma$ values prevent the low-pass filter from effectively extracting the primary representation of the KV matrix (Base KV). This failure to identify the Outlier KV consequently leads to a decline in model performance. This phenomenon is consistent with our core findings.
@@
Method H2O [49] SnapKV [21] LOOK-M [35] MEDA [34] FlashCache